OpenStudy (anonymous):

Using graphing technology, create your own quadratic equation whose graph opens up and does not cross the x-axis. Use complete sentences to explain what piece of the quadratic equation causes its graph to open up or down. Explain what the graph indicates about the type of solutions for the quadratic equation.....I am not sure how to do this problem ??

OpenStudy (shayaan_mustafa):

Hi 2523 :) It is the question about parabola.

OpenStudy (anonymous):

Yes :)

OpenStudy (shayaan_mustafa):

\[\large y=4px^2\] It is the parabola which opens upward having center at origin.

OpenStudy (anonymous):

Huh ? It is supossed to open upwards and not cross the x-axis

OpenStudy (shayaan_mustafa):

At origin, it will just touches x-axis but not cross x-axis. And it is the focus which causes it to open up and down. \[\large y=-4px^2\] It is the parabola which opens downward P=focus of the parabola.

OpenStudy (anonymous):

So are you saying that is the equation i am supoosed to put in ? and also i need to find the vertex . How am i supossed to find the vertex of what you gave me?

OpenStudy (anonymous):

Would that open up ?

OpenStudy (shayaan_mustafa):

Wait please. Until I close another question then I reply you. Don't go away. OK. :)

OpenStudy (shayaan_mustafa):

Well continue.

OpenStudy (anonymous):

Okay

OpenStudy (shayaan_mustafa):

Of course it is the required equation. \[\large y=4px^2\] And it opens upward. For Vertex: Let have a look at general equation of parabola. \[\large (y-k)=4p(x-h)^2\] where (h,k) defines its vertex. If (h,k)=(0,0) then the general equation of parabola reduces to standard equation i.e. \[\large (y-0)=4p(x-0)^2\] \[\large y=4px^2\] Got it uptill now?

OpenStudy (anonymous):

i need a quadratic equation in standard for ie y=ax^2 +bx+c

OpenStudy (anonymous):

that doesnt cross x axis and opens upwards when solved

OpenStudy (shayaan_mustafa):

@2523 \[\Large y=4px^2\]is the quadratic equation. If you solve this then this give you two answers . And also opens upward.